OpenStudy (anonymous):

The position of an object at time t is given by s(t) = -2 - 6t. Find the instantaneous velocity at t = 2 by finding the derivative.

OpenStudy (anonymous):

@hartnn @paki

hartnn (hartnn):

have you learnt derivatives ?

OpenStudy (anonymous):

yes

hartnn (hartnn):

so could you find the derivative of s(t) ?

OpenStudy (anonymous):

would that be \[\frac{ -2-6(t+h)-2-6(t) }{ ? }\]

OpenStudy (anonymous):

?=h

hartnn (hartnn):

thats by definition, haven't you taught direct formulas ?

hartnn (hartnn):

like \(\Large \dfrac{d}{dx} (x^n) = n x^{n-1}\)

hartnn (hartnn):

and derivative of a constant is simply 0

OpenStudy (anonymous):

so how would i right thaat out with the numbers?

hartnn (hartnn):

 see if that makes sense ^^